# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ 4 Discussion
Our proposed LambdaG approach matches many desiderata of an AV method:
  1. 1.
    It works on both short and long documents;
  2. 2.
    It is not affected or minimally affected by topic or content;
  3. 3.
    It is stable to small genre variations in the reference corpus;
  4. 4.
    It is more interpretable by an analyst;
  5. 5.
    It is not computationally intensive and does not have a high run time;
  6. 6.
[…]
  7. 7.
    It has a theoretical foundation based on Cognitive Linguistic theories of language processing.
Section 3 showed evidence of the first six of these points as well as evidence that LambdaG outperforms other AV state-of-the-art methods. We believe that a possible reason to explain these results, especially considered the simplicity of the method compared to approaches such as SiamBERT, is precisely that LambdaG has a number of properties that make it consistent with knowledge about language processing gained from studies in Cognitive Science and Cognitive Linguistics.
@@
Since the pioneering study of Mosteller and Wallace [79], it is well known that the frequency of functional items distinguish authors and this conclusion is also evident in the case of LambdaG. The reason why this is the case, however, has always been seen as a mystery in the field of authorship analysis and stylometry [56]. A connection between the unreasonable effectiveness of these functional items and Cognitive Linguistics has been proposed by Nini [82] and we believe that the success of LambdaG is additional evidence for these explanations.
Cognitive Linguistics models language as a complex adaptive system [9], where regularities emerge from the complexity of language used in interactions, like a phenomenon of the third kind [33]. This is partially evidenced by decades of linguistics studies on corpora which have revealed that the traditional conception of a grammar as an abstract set of phrase structure rules is untenable and that, instead, a better model of grammar is a lexicogrammar continuum that goes from fixed lexicalized expressions (e. g., “I don’t know”) to flexible templates (e. g., “the” X-er “the” Y-er) up to completely abstract structures (e. g., Doer Action Object).
This continuum roughly corresponds to the continuum between declarative memory, which is semantic/encyclopedic memory for facts and which tends to be conscious, and procedural memory, which is instead memory for the probabilistic prediction of patterns of sequences and that deals with subconscious knowledge of habits and skills [108]. The roles that these two memories play in language is clearly evidenced in cases of aphasia. When an impairment affects circuits of declarative memory, this gives rise to fluent aphasia, resulting in speech that is mostly constituted by function words like articles, prepositions, pronouns, and morphological forms but with lack of content words that contain meaning, leading to fluent but nonsensical speech comparable to POSNoise pre-processed data such as the one in Table 5. Instead, when the impairment is on procedural memory, then speech is only consisting of content words that are not connected through grammar [109]. These two memory systems work together and, typically, once a habit is formed through repetition then procedural memory takes over, leading to fluency of expression. This transition explains why languages tend to be heavily formulaic. For example, in English one can say big mistake but not large mistake, even though the formal grammar would seemingly have no problems with the latter expression, which is an adjective modifying a noun. What explains this reliance on formulaic expressions is the cognitive tendency to recur to memory as much as possible rather than creating something new every time we use language [108].
The way these units of procedural memory are formed is through the process of chunking: when two items that often go together are found together, then these two items can be treated as one, thus saving space in working memory [39, 21, 28]. In all effects, chunking is a process of information compression and, if the units of grammar are chunks, then the logical inference is that grammar itself can be thought of as a compressed version of a language or, equivalently, as a probability model. A probability model of only functional items is therefore quite like the Grammar Models used in this study.
Cognitive Linguistics also explains why each individual has a different and possibly unique grammar, a statement encapsulated by the Principle of Linguistic Individuality [82]. The status of unithood given by chunking is not discrete (e. g., something is or is not a unit) but a continuum of strength of unithood, where this strength is called entrenchment [32, 97]. Because entrenchment is dependent on exposure and usage, it is fundamentally personal and idiosyncratic. It is for this reason that Langacker hypothesized that “the set of units mastered by all members of a speech community might be a rather small proportion of the units constituting the linguistic ability of any given speaker” [71], a hypothesis that is finding increasing empirical support [82]. For example, for most if not all speakers of English the entrenchment for a sequence such as “of the” is maximal, meaning that they can produce it effortlessly and unconsciously as a single unit. In contrast, a sequence such as “the of” would not be a unit for speakers of English because it is virtually never encountered or used. If a speaker of English encounters this sequence, they would decompose it as a structure made up of two units, “the” and “of”. In between these two extreme examples there is a lot of freedom as to what counts as a unit for a person or another, depending on what each individual has automatized in their procedural memory. For example, an individual $\mathcal{A}_{1}$ might have a very entrenched sequence such as “it is known that” when writing academic articles while an individual $\mathcal{A}_{2}$ might have “it is well established that”. These sequences will then have open slots ready to be filled in with the appropriate meaning. This kind of phenomenon theorized by Langacker can be seen at play in this method, especially when looking at the examples of idiosyncratic structures in Section 6. As briefly mentioned in Section 5.3, therefore, the likelihood of a token given a context assigned by an author $\mathcal{A}$’s Grammar Model, $P(t|c;G_{\mathcal{A}})$, can be seen as a way to measure the level of entrenchment of token $t$ in context $c$ for $\mathcal{A}$ The higher the entrenchment, the more likely it is that $\mathcal{A}$ is using procedural memory and that it is therefore performing an unconscious routine that constitutes a unit for this individual.
There have been other attempts at modeling the grammar of an author in the past for AV or AA purposes [24, 7, 37] but these methods relied too much on Part of Speech tags or on syntactic trees which are often not compatible with predictions made in Cognitive Science and Cognitive Linguistics. As explained above, the picture of grammar painted by Cognitive Linguistics is much more lexically-based and word-idiosyncratic. For example, the preposition “of” in English behaves very oddly compared to other prepositions [100]. Giving it the same Part of Speech label as to other prepositions is therefore going to obscure a lot of important grammatical information. Instead, a topic-masking method such as POSNoise is much more consistent with the predictions of Cognitive Linguistics because it leaves the function items unaffected while only obscuring those items are more likely to change from context to context depending on the communicative needs.
@@
To sum up, the model of an individual’s grammar that emerges from Cognitive Science and Cognitive Linguistics is one that is very much compatible with the Grammar Models used for LambdaG. Although language modeling is seen in computational linguistics as a tool and not necessarily a realistic representation of language, evidence from Cognitive Linguistics suggests that, in reality, a language model, while not being realistic as a model of linguistic production, is instead a plausible representation of how knowledge of grammar is stored in the mind. The mechanisms of such a sequential processing can be efficiently represented through the short-term correlations captured by $n$-gram models as done in this work.
@@
A deeper connection between LambdaG and approaches to AV based on language modeling or, as for COAV, on compression algorithms can also be drawn, thus offering a plausible explanation for their success. Through the lenses of Information Theory, the average $P(s;G)$ for each $s\in\mathcal{D}$ can be also seen as an estimate of the cross-entropy rate of the Grammar Model $G$ against the probability distribution of grammatical tokens that generates $\mathcal{D}$. In language modeling literature, the cross-entropy is often exponentiated to calculate the perplexity of a document given a model. Because for LambdaG the two alternative Grammar Models are compared in relation to the same $\mathcal{D}$, for the purposes of this AV application, $\lambda_{G}$ is equivalent to comparing the perplexities of the candidate author’s model vs. the reference model given the same $\mathcal{D}$ and assign the document to the candidate author if this perplexity is lower. The cross-entropy rate can also be interpreted as the average length in bits of a binary encoding of grammatical tokens for a data-compression scheme that is optimized for the probability distribution $P(\cdot;G)$. For this reason, a language model can be seen as a text-compression scheme. This connection thus links compression and language model AV methods like LambdaG and COAV to the Cognitive Linguistics conception of an individual’s grammar as a compressed representation of real language usage encountered by said individual.
@@
In addition to the successful results of LambdaG, we note that the qualitative explorations reported in Section 6 also constitute evidence that LambdaG captures linguistic behavior compatible with these Cognitive Linguistic explanations. Although it is impossible to know for certain the effective mental status of the sequences found in the examples in Section 6, we propose that it is plausible to believe that the authors of these texts do not know that they use the sequences of tokens highlighted in red or that these sequences identify them in the general population. These sequences, which are often seemingly unremarkable (e. g.,enough to be ADVERB VERB” or “so PRONOUN cant”), can be extremely idiosyncratic and uniquely used only by the candidate author. Given the fact that the problems analyzed in this paper involve relatively short texts (the longest texts are just around 2000 tokens), we must conclude that these idiosyncratic units are easy to find and this conclusion therefore strongly supports Langacker’s proposal that most units in a language are not shared by many individuals.
@@
We believe that additional evidence towards this assertion is given by the results of the cross-reference experiments. Controlling for genre compatibility is probably the hardest problem in AV and AA. LambdaG is very robust to this variation in many cases, much more than expected given the results of similar tests carried out on IM [41]. This is despite the incredible challenges in some of the comparisons, such as in the analysis of the $\mathcal{C}_{\mathrm{PJ}}$ chat logs using as reference the academic papers in $\mathcal{C}_{\mathrm{ACL}}$. The genre differences between these two corpora cannot be greater and therefore it is hard to believe that the method is not performing at chance level. A possible explanation for the success of LambdaG in these circumstances is that the reference corpus effectively does not play a major role. If individual grammars are so different from each other, then the role of the reference corpus could be to just weight down the those units that are well-entrenched for most individuals (e. g., of the) and that therefore carry almost no authorial information. If this explanation is correct, then this means that once these common units are taken into account, the rest is idiosyncratic enough to be identifying.
@@
Finally, due to the fact that these idiosyncratic units are easy to find within as little as 1,000 to 2,000 tokens, we add that a consequences of the present findings and of the underlying Cognitive Linguistic theory is that the grammar of a person can be seen as a behavioral biometric [82]. Similarly to handwriting, gait, typing patterns, computer mouse usage, and most features of human voice, grammar is also a cognitive process relying on procedural memory that strengthens with habit and repetition. We propose that this is the explanation underlying most if not all AV and AA methods that, by measuring the frequency of functional items, are indirectly analyzing the idiosyncratic grammar of an individual. LambdaG tends to be more successful probably because it is measuring the same phenomenon more directly.
@@ 4.1 Limitations and future research
We argue that future research should therefore continue in this direction of extracting better representations of the grammar of an individual. The POSNoise algorithm used here is language dependent and suffers from the use of general part of speech categories that are not compatible with Cognitive Linguistics, which instead predicts the existence of categories that are more semantic in nature and at lower degrees of abstractness. New future methods should embrace the close relationship between Cognitive Linguistics and Information Theory to gain more insight into grammar, which could consequently lead to further breakthroughs in AV.
In addition, LambdaG is still affected by genre differences and performs less well in registers where there is less variation, such as academic prose. These two last limitations affect more or less all existing algorithms which are outperformed by LambdaG, which could mean that these limitations are intrinsic to AV and, indeed, there would be good linguistic reasons for why this would be the case. Despite these limitations, future methods based on a better understanding of grammar could still lead to more progress. In this respect, the reliance on a reference corpus could also be seen as a limitation of any binary AV method. The tendency for future AV methods should be to increase the power of unary methods that only need a test dataset.
Finally, another limitation of the present paper is that the research was carried out entirely on English. The theory underlying LambdaG generalizes to other languages and for this reason the prediction is that it would also outperform the same baseline methods in other languages. The way Grammar Models are trained in languages with a rich morphology is likely to be different, however. Different tokenizers might have to be employed as well as different algorithms such as POSNoise to focus the model on the grammatical aspects of the text. Another goal for the future would therefore be to find a language-independent variant of LambdaG.
@@ 4.2 Final Remarks
The advances in AV made in this paper were possible because of the collaboration between Linguistics and Natural Language Processing. In the 2022 edition of the PAN competition on AV, Stamatatos et al. [104] concluded that “all submissions, despite their increased level of sophistication in most of the cases, were outperformed by a naive baseline based on character $n$-grams and cosine similarity” thus noticing a situation of stall in the progress on AV. Knowledge from Linguistics could resolve these stalls, as demonstrated in this paper, by shifting the point of view rather than adding more computational sophistication or making algorithms more complex. The method and philosophy introduced in the present work offer significant advancements in integrating Linguistics theory and Computer Science methods towards a new paradigm in AA and AV where these two fields are better harmonized. The task of identifying a writer from the language used in a text is really an application of Linguistics and there are various reasons for computational methods to maintain a close relationship with Linguistics. Firstly, these methods directly or indirectly exploit linguistic patterns that require knowledge of Linguistics to be described and understood. Secondly, knowledge of Linguistics is needed to understand and potentially explain in a court of law the reasons why these patterns identify or exclude a certain author. We believe that the work in this paper has made some significant advances in AV while also proposing and stimulating a paradigm change to move on from the impasse described by Stamatatos [104] towards new AV methods that are not only well-performing but also scientifically meaningful.
@@ 5 Methods
In this section, we present the more technical details of our work, first with an introduction to the corpora we have compiled to evaluate LambdaG, then by providing more details about the POSNoise content-masking algorithm as well as the details of the Kneser-Ney smoothing algorithm, the likelihood ratio framework for forensic sciences and, finally, the Siamese neural network baseline method, SiamBERT, developed for this analysis.
@@ 5.1 Corpora
To evaluate LambdaG, we compiled a collection of 12 corpora covering a variety of challenges. These include, for instance, very short texts ($\approx 5$ sentences), texts with a lack of coherence, texts with excessive use of slang, texts (of the same author) written at different times, texts with opposing topics, texts with distorted writing style caused by post-editing, texts with truncated sentences as well as non-standard spellings and missing spaces. In total, the corpora comprise 17,306 verification cases, were each corpus $\mathcal{C}=\{c_{1},c_{2},\ldots\}$ is split into author-disjunct training and test partitions based on a 40/60 ratio, resulting in 6,402 training and 10,904 test verification cases. Each $c\in\mathcal{C}$ denotes a verification case $(\mathcal{D}_{\mathcal{U}},\mathbb{D}_{\mathcal{A}})$, where $\mathcal{D}_{\mathcal{U}}$ is an unknown document and $\mathbb{D}_{\mathcal{A}}$ is a set of example documents of the known author $\mathcal{A}$. To counteract population homogeneity, a known bias in authorship verification described by Bevendorff [11], we ensured that for each author $\mathcal{A}$ there are exactly two (one same-authorship (Y) and one different-authorship (N)) verification cases. Furthermore, we constructed all corpora so that the number of Y- and N-cases is equal, in other words, all training and test corpora are (strictly) balanced. In what follows, we describe all corpora in detail, underline their respective challenges and explain the pre-processing [^2] The pre-processing of the documents was carried out using TextUnitLib https://github.com/Halvani/TextUnitLib steps we have performed. A compact overview of the 12 corpora, including their partitioning into training and test corpora, as well as further information can be found in Table 4.
@@
Table 4: All used corpora and their key statistics. Notation: $|\mathcal{C}|$ denotes the number of verification cases in each corpus $\mathcal{C}$, while avg$(\mathbb{D}_{\mathcal{A}})$ and avg$(\mathcal{D}_{\mathcal{U}})$ denotes the average token count (and character count in parentheses) of all $\mathcal{D}_{\mathcal{A}}\in\mathbb{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$, respectively.

### v2

@@ 5 Discussion
We believe that our proposed LambdaG approach matches many desiderata of an AV method. The evaluation on our set of 12 corpora demonstrates that LambdaG works consistently well for relatively short texts in a variety of different genres. In addition, because the method only considers information in functional items, LambdaG is also minimally affected by potential content biases in the data, such as named entities or topic. In contrast to the other very successful binary-extrinsic AV method, IM, LambdaG tends to be much more robust to variations in the reference corpus. Moreover, unlike IM, which relies on the similarity between vectors of thousands of character $n$-grams, LambdaG generates text heat maps that aid the analyst’s understanding of the output of the method. Finally, LambdaG is unique among other AV methods in being motivated by Cognitive Linguistics and in having a theoretical explanation for its mechanisms.
@@
We believe that its greater scientific plausibility justifies LambdaG over other state-of-the-art methods. Since the pioneering study of Mosteller and Wallace [94], it is well known that the frequency of functional items distinguish authors and this conclusion is also evident in our results. The reason why this is the case, however, has always been seen as a mystery in the field of authorship analysis and stylometry [69]. A connection between the unreasonable effectiveness of these functional items and Cognitive Linguistics has been proposed by Nini [97] and we believe that the success of LambdaG is additional evidence for these explanations.
@@
As outlined in Section 2, the mental representation of the grammar for an individual can be seen as a a probability model over lexicogrammatical chunks quite like the Grammar Models used in LambdaG. Although language modeling is seen in computational linguistics as a tool and not necessarily a realistic representation of language, evidence from Cognitive Linguistics suggests that, in reality, a language model, while not being realistic as a model of linguistic production, is instead a plausible representation of how knowledge of grammar is stored in the mind, as well as an in silico model of the brain language system [127]. The mechanisms of such a sequential processing can be efficiently represented through the short-term correlations captured by $n$-gram models as done in this work.
@@
A deeper connection between LambdaG and approaches to AV based on language modeling or, as for COAV, on compression algorithms can also be drawn, thus offering a plausible explanation for their success. Through the lenses of Information Theory, the average $P(S;G)$ for each sentence $S$ in $\mathcal{D}$ can be also seen as an estimate of the cross-entropy rate of the Grammar Model $G$ against the probability distribution of grammatical tokens that generates $\mathcal{D}$. In language modeling literature, the cross-entropy is often exponentiated to calculate the perplexity of a document given a model. Because for LambdaG the two alternative Grammar Models are compared in relation to the same $\mathcal{D}$, for the purposes of this AV application, $\lambda_{G}$ is equivalent to comparing the perplexities of the candidate author’s model vs. the reference model given the same $\mathcal{D}$ and assign the document to the candidate author if this perplexity is lower. The cross-entropy rate can also be interpreted as the average length in bits of a binary encoding of grammatical tokens for a data-compression scheme that is optimized for the probability distribution $P(\cdot;G)$. For this reason, a language model can be seen as a text-compression scheme. This connection thus links compression and language model AV methods like LambdaG and COAV to the Cognitive Linguistics conception of an individual’s grammar as a compressed representation of real language usage encountered by said individual.
@@
In addition to the successful results of LambdaG, we note that the qualitative explorations reported in Section 7 also constitute evidence that LambdaG captures linguistic behavior compatible with these Cognitive Linguistic explanations. That individuals vary greatly in their mental grammar, as theorized by Langacker, can be seen at play in these text heat maps. The theoretical interpretation of these results is that a stronger shade of red effectively means a greater likelihood that the token is entrenched for that individual’s unique grammar in contrast to the population. Although it is impossible to know for certain the effective mental status of the sequences found in the examples in Section 7, we propose that it is plausible to believe that the authors of these texts do not know that they use the sequences of tokens highlighted in red or that these sequences identify them in the general population. These sequences, which are often seemingly unremarkable (e. g.,enough to be ADVERB VERB” or “so PRONOUN cant”[sic]), can be extremely idiosyncratic and uniquely used only by the candidate author. Given the fact that the problems analyzed in this paper involve relatively short texts (the longest texts are just around 2000 tokens), we must conclude that these idiosyncratic units are easy to find and this conclusion therefore strongly supports Langacker’s proposal that most units in a language are not shared by many individuals.
@@
We believe that additional evidence towards this assertion is given by the results of the cross-reference experiments. Controlling for genre compatibility is probably the hardest problem in AV and AA. LambdaG is very robust to this variation in many cases, much more than expected given the results of similar tests carried out on IM [51]. This is despite the incredible challenges in some of the comparisons, such as in the analysis of the $\mathcal{C}_{\mathrm{PJ}}$ chat logs using as reference the academic papers in $\mathcal{C}_{\mathrm{ACL}}$. The genre differences between these two corpora cannot be greater and therefore it is hard to believe that the method is not performing at chance level. A possible explanation for the success of LambdaG in these circumstances is that the reference corpus effectively does not play a major role. If individual grammars are so different from each other, then the role of the reference corpus could be to just weight down the those units that are well-entrenched for most individuals (e. g.,of the”) and that therefore carry almost no authorial information. If this explanation is correct, then this means that once these common units are taken into account, the rest is idiosyncratic enough to be identifying.
@@
Finally, due to the fact that these idiosyncratic units are easy to find within as little as 1,000 to 2,000 tokens, we add that a consequences of the present findings and of the underlying Cognitive Linguistic theory is that the grammar of a person can be seen as a behavioral biometric [97]. Similarly to handwriting, gait, typing patterns, computer mouse usage, and most features of human voice, grammar is also a cognitive process relying on procedural memory that strengthens with habit and repetition. We propose that this is the explanation underlying most if not all AV and AA methods that, by measuring the frequency of functional items, are indirectly analyzing the idiosyncratic grammar of an individual. It is possible that LambdaG is more successful because it is measuring this same phenomenon more directly.
@@ 5.1 Limitations and Future Research
We argue that future research should therefore continue in this direction of extracting better representations of the grammar of an individual. The POSNoise algorithm used here is language dependent and suffers from the use of general part of speech categories that are actually not compatible with Cognitive Linguistics, which instead predicts the existence of categories that are more semantic in nature and at lower degrees of abstraction. New future methods should embrace the close relationship between Cognitive Linguistics and Information Theory to gain more insight into grammar, which could consequently lead to further breakthroughs in AV.
In addition, LambdaG is still affected by genre differences and performs less well in registers where there is less variation, such as academic prose. These two last limitations affect more or less all existing algorithms which are outperformed by LambdaG, which could mean that these limitations are intrinsic to AV and, indeed, there would be good linguistic reasons for why this would be the case. Despite these limitations, future methods based on a better understanding of grammar could still lead to more progress. In this respect, the reliance on a reference corpus could also be seen as a limitation of any binary-extrinsic AV method. The tendency for future AV methods should be to increase the power of unary methods that consider only the documents within a verification case.
LambdaG also suffers from two limitations that also apply to IM because of their bootstrapping component: both algorithms are stochastic and a higher run time due to the number of repetitions. In our experiment we noted that LambdaG is faster than IM, although this depends on the specific implementation. Our test of the $r$ hyperparameter also shows that 30 repetitions are enough for best performance, in contrast to the 100 iterations in IM. In addition, the computational complexity of LambdaG is still very limited compared to any method using neural networks and running on GPUs. To mitigate the effect of randomness in the bootstrapping, all experiments in this paper were run five times. However, a deterministic algorithm would definitely be preferable.
Finally, another limitation of the present paper is that the research was carried out entirely on English. The theory underlying LambdaG generalizes to other languages and for this reason the prediction is that it would also outperform the same baseline methods in other languages. The way Grammar Models are trained in languages with a rich morphology is likely to be different, however. Different tokenizers might have to be employed as well as different algorithms such as POSNoise to focus the model on the grammatical aspects of the text. Our test showing that POS tags are not necessary is also encouraging for the replicability of this algorithm for languages with less computational resources. In general, though, another goal for the future would therefore be to find a language-independent variant of LambdaG.
@@ 5.2 Final Remarks
The advances in AV made in this paper were possible because of the collaboration between Linguistics and Natural Language Processing. In the 2022 edition of the PAN competition on AV, Stamatatos et al. [121] concluded that “all submissions, despite their increased level of sophistication in most of the cases, were outperformed by a naive baseline based on character $n$-grams and cosine similarity” thus noticing a situation of stall in the progress on AV. Knowledge from Linguistics could resolve these stalls, as demonstrated in this paper, by shifting the point of view rather than adding more computational sophistication or making algorithms more complex. The method and philosophy introduced in the present work offer significant advancements in integrating Linguistics theory and Computer Science methods towards a new paradigm in AA and AV where these two fields are better harmonized. The task of identifying a writer from the language used in a text is really an application of Linguistics and there are various reasons for computational methods to maintain a close relationship with Linguistics. Firstly, these methods directly or indirectly exploit linguistic patterns that require knowledge of Linguistics to be described and understood. Secondly, knowledge of Linguistics is needed to understand and potentially explain in a court of law the reasons why these patterns identify or exclude a certain author. We believe that the work in this paper has made some significant advances in AV while also proposing and stimulating a paradigm change to move on from the impasse described by Stamatatos [121] towards new AV methods that are not only well-performing but also scientifically plausible.
@@ 6 Methods
In this section, we present the more technical details of our work, first with an introduction to the corpora we have compiled to evaluate LambdaG, then by providing more details about the POSNoise content-masking algorithm as well as the details of the Kneser-Ney smoothing algorithm, the likelihood ratio framework for forensic sciences and, finally, our own baseline method, SiamBERT, developed for this analysis.
@@ 6.1 Corpora
To evaluate LambdaG, we compiled a collection of 12 corpora covering a variety of challenges. These include, for instance, very short texts ($\approx 5$ sentences), texts with a lack of coherence, texts with excessive use of slang, texts (of the same author) written at different times, texts with opposing topics, texts with distorted writing style caused by post-editing, texts with truncated sentences as well as non-standard spellings and missing spaces. In total, the corpora comprise 17,306 verification cases, were each corpus $\mathcal{C}=\{c_{1},c_{2},\ldots\}$ is split into author-disjunct training and test partitions based on a 40/60 ratio, resulting in 6,402 training and 10,904 test verification cases. Each $c\in\mathcal{C}$ denotes a verification case $(\mathcal{D}_{\mathcal{U}},\mathbb{D}_{\mathcal{A}})$, where $\mathcal{D}_{\mathcal{U}}$ is an unknown document and $\mathbb{D}_{\mathcal{A}}$ is a set of example documents of the known author $\mathcal{A}$. To counteract population homogeneity, a known bias in AV described by Bevendorff [14], we ensured that for each author $\mathcal{A}$ there are exactly two verification cases i. e., one with same-authorship (Y) and one with different-authorship (N). Furthermore, we constructed all corpora so that the number of Y- and N-cases is equal, in other words, all training and test corpora are (strictly) balanced. In what follows, we describe all corpora in detail, underline their respective challenges and explain the respective pre-processing steps (using TextUnitLib [52]) we have performed for each corpus. A compact overview of the 12 corpora, including their partitioning into training and test corpora, as well as further information can be found in Table 5.
@@
Table 5: All corpora used and their key statistics. Notation: $|\mathcal{C}|$ denotes the number of verification cases in each corpus $\mathcal{C}$, while avg$(\mathbb{D}_{\mathcal{A}})$ and avg$(\mathcal{D}_{\mathcal{U}})$ denotes the average token count (and character count in parentheses) of all $\mathcal{D}_{\mathcal{A}}\in\mathbb{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$, respectively.